# Gaussian Scenes: Pose-Free Sparse-View Scene Reconstruction using Depth-Enhanced Diffusion Priors

### v1

@@ Abstract
In this work, we introduce a generative approach for pose-free reconstruction of $360^{\circ}$ scenes from a limited number of uncalibrated 2D images. Pose-free scene reconstruction from incomplete, unposed observations is usually regularized with depth estimation or 3D foundational priors. While recent advances have enabled sparse-view reconstruction of unbounded scenes with known camera poses using diffusion priors, these methods rely on explicit camera embeddings for extrapolating unobserved regions. This reliance limits their application in pose-free settings, where view-specific data is only implicitly available. To address this, we propose an instruction-following RGBD diffusion model designed to inpaint missing details and remove artifacts in novel view renders and depth maps of a 3D scene. We also propose a novel confidence measure for Gaussian representations to allow for better detection of these artifacts. By progressively integrating these novel views in a Gaussian-SLAM-inspired process, we achieve a multi-view-consistent Gaussian representation. Evaluations on the MipNeRF360 dataset demonstrate that our method surpasses existing pose-free techniques and performs competitively with state-of-the-art posed reconstruction methods in complex $360^{\circ}$ scenes.
@@ 1 Introduction
Despite several previous attempts (Deng et al., 2022; Roessle et al., 2022; Li et al., 2024; Xiong et al., 2023; Zhu et al., 2023; Chung et al., 2023; Wang et al., 2023a; Jain et al., 2021; Niemeyer et al., 2022; Wynn and Turmukhambetov, 2023) in the field of sparse novel view synthesis, only a few - CF-3DGS (Fu et al., 2023), COGS Jiang et al. (2024), InstantSplat (Fan et al., 2024) have been successful in dealing with the problem without precomputed camera parameters. These methods jointly optimize Gaussians and camera parameters during training and use different 2D priors for regularization. Although they find success in artifact and blur reduction, they are less effective in $360^{\circ}$ scene reconstruction due to a lack of generative priors.
Reconstructing full $360^{\circ}$ scenes from limited views requires robust priors, such as those from powerful 2D image or video diffusion models (Nichol et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022) that encode common 3D structures. Recent methods (Liu et al., 2023; Sargent et al., 2024; Wu et al., 2024; Gao* et al., 2024; Blattmann et al., 2023) incorporate view and context conditioning into these models, fine-tuning them on extensive real-world and synthetic multi-view datasets to support radiance field optimization. This enables realistic extrapolation into unobserved areas of complex $360^{\circ}$ scenes. However, these models depend on accurate pose information, making them unsuitable for scenarios where pose estimates deviate significantly from ground truth, leading to misaligned view generation. To our knowledge, only Gaussian Object (Yang et al., 2024), iFusion (Wu et al., 2023), and UpFusion (Nagoor Kani et al., 2024) provide pose-free generative solutions, but they are designed specifically for 3D object reconstruction, not large-scale scenes.
In the absence of camera conditioning, we augment Stable Diffusion Rombach et al. (2022) model with additional channels for context, 3DGS render, and depth maps with Gaussian artifacts, and a pixel-aligned confidence map capturing missing regions and reconstruction artifacts in the RGBD image. During inference, the model predicts a clean, inpainted version of the conditioning artifact image and an aligned depth map. This formulation prevents a pose-conditioning requirement common with most multiview diffusion models while also using limited multiview data for finetuning. This also alleviates dependency on large-scale 3D datasets which are usually synthetic and low quality compared to real-world scenes.
@@
We present GScenes, an efficient method that uses 3D foundational (Leroy et al., 2024) and RGBD diffusion priors for pose-free sparse-view reconstruction of complex $360^{\circ}$ scenes. We first estimate a point cloud and approximate scene intrinsics and extrinsics using DUSt3R and then jointly optimize both Gaussians and camera parameters with just a few iterations of 3DGS. We then generate novel views by sampling from a B-spline trajectory fitted to the training views. Resulting renders and depth maps contain missing details and Gaussian artifacts that are refined by our diffusion prior and then used to further optimize the 3D Gaussians and camera poses. Despite using weaker generative priors compared to state-of-the-art pose-dependent methods, we show competitive qualitative and quantitative performance with ReconFusion (Wu et al., 2024) and CAT3D (Gao* et al., 2024) and comprehensively outperform all other pose-dependent and pose-free techniques on reconstruction quality. Our method leverages stronger priors than simple regularizers while not relying on million-scale multi-view data or huge compute resources to train a 3D-aware diffusion model.
In summary, our contributions are as follows.
  • We propose a confidence measure to effectively detect empty regions, floaters, and artifacts in novel view renders and depth maps from 3DGS.
  • We present an instruction-following RGBD diffusion model capable of synthesizing plausible novel views in a 3DGS scene initialized from sparse unposed images.
  • We show that integrating our diffusion priors with a strong geometry prior in DUSt3r facilitates fast reconstruction of a 3D scene previously only possible with pose-dependent finetuned video diffusion priors.
  • We show that GScenes outperforms recent works based on regularization and generative priors in the reconstruction of large 3D scenes.
2 Related Work
2.1 Sparse-View Radiance Fields
After NeRF Mildenhall et al. (2020) and 3D Gaussian Splatting Kerbl et al. (2023) advanced inverse rendering of radiance fields, numerous methods emerged to reduce dense scene capture needs to sparse views, categorized into regularization techniques and generalizable reconstruction priors.
Regularization Techniques
Fitting a 3D representation from sparse observations is an ill-posed problem prone to local minima, often manifesting as ’floaters’ in novel view renderings of radiance fields. Regularization techniques are commonly used to address limited data. Existing methods leverage depth from Structure-from-Motion (SfM)Deng et al. (2022); Roessle et al. (2022), monocular estimationLi et al. (2024); Xiong et al. (2023); Zhu et al. (2023); Chung et al. (2023), or RGB-D sensors Wang et al. (2023a). DietNeRF Jain et al. (2021) uses a semantic consistency loss based on CLIP Radford et al. (2021) features, while FreeNeRF Yang et al. (2023) regularizes the frequency range of NeRF inputs. Approaching generative priors, RegNeRF Niemeyer et al. (2022) and DiffusioNeRF Wynn and Turmukhambetov (2023) maximize the likelihoods of rendered patches using normalizing flows or diffusion models, respectively.
Generalizable Reconstruction
When only a few or a single view is available, regularization techniques are often insufficient to resolve reconstruction ambiguities. To address this, recent research focuses on training priors for novel view synthesis across multiple scenes. pixelNeRF Yu et al. (2021) uses pixel-aligned CNN features as conditioning for a shared NeRF MLP, while other approaches Trevithick and Yang (2021); Chen et al. (2021); Henzler et al. (2021); Lin et al. (2023b) condition NeRF on 2D or fused 3D features. Further priors have been learned on triplanes Irshad et al. (2023), voxel grids Guo et al. (2022), and neural points Wewer et al. (2023). Leveraging 3D Gaussian Splatting Kerbl et al. (2023), methods like pixelSplat Charatan et al. (2024) and MVSplat Chen et al. (2024) achieve state-of-the-art performance in stereo view interpolation and real-time rendering. However, these regression-based models often produce blurry outputs under high uncertainty. In contrast, generative methods like GeNVS Chan et al. (2023) and latentSplat Wewer et al. (2024) aim to sample from multi-modal distributions, offering better handling of ambiguous novel views.
Generative Priors
Diffusion Models
In recent years, diffusion models Ho et al. (2020); Dhariwal and Nichol (2021) emerged as the state-of-the-art for image synthesis. They are characterized by a pre-defined forward noising process that gradually destroys data by adding random (typically Gaussian) noise. The objective is to learn a reverse denoising process with a neural network that, after training, can sample from the data distribution given pure noise. Important improvements include refined sampling procedures Song et al. (2021); Karras et al. (2022) and the more efficient application in a spatially compressed latent space compared to the high-resolution pixel space Rombach et al. (2022). Their stable optimization, in contrast to GANs, enabled today’s text-to-image generators Nichol et al. (2022); Ramesh et al. (2022); Saharia et al. (2022); Rombach et al. (2022) trained on billions of images Schuhmann et al. (2022). In recent years, diffusion models Ho et al. (2020); Dhariwal and Nichol (2021) have become state-of-the-art for image synthesis. These models use a forward noising process to progressively add random (typically Gaussian) noise, and train a neural network to reverse this process, enabling sampling from the data distribution given pure noise. Key improvements include refined sampling techniques Song et al. (2021); Karras et al. (2022) and more efficient application in compressed latent spaces rather than high-resolution pixel spaces Rombach et al. (2022). Their stable optimization, unlike GANs, has enabled the success of text-to-image generators Nichol et al. (2022); Ramesh et al. (2022); Saharia et al. (2022); Rombach et al. (2022) trained on billions of images Schuhmann et al. (2022).
2D Diffusion for 3D While diffusion models have been applied directly on 3D representations like triplanes Shue et al. (2023); Chen et al. (2023a), voxel grids Müller et al. (2023), or (neural) point clouds Zhou et al. (2021); Melas-Kyriazi et al. (2023); Schröppel et al. (2024), 3D data is scarce. Given the success of large-scale diffusion models for image synthesis, there is a great research interest in leveraging them as priors for 3D reconstruction and generation. DreamFusion Poole et al. (2023) and follow-ups Wang et al. (2023b); Lin et al. (2023a); Chen et al. (2023b); Deng et al. (2023); Tang et al. (2024) employ score distillation sampling (SDS) to iteratively maximize the likelihood of radiance field renderings under a conditional 2D diffusion prior. For sparse-view reconstruction, existing approaches incorporate view-conditioning via epipolar feature transform Zhou and Tulsiani (2023), cross-attention to encoded relative poses Liu et al. (2023); Sargent et al. (2024), or pixelNeRF Yu et al. (2021) feature renderings Wu et al. (2024). However, this fine-tuning is expensive and requires large-scale multi-view data, which we circumvent with GScenes.
@@ 2.1 Sparse-View Radiance Fields
3 Method
This section begins with an overview of our pipeline in Sec 3.1. In Sec 3.3, we present a confidence measure for 3DGS rasterization to capture artifacts and missing regions in rendered images and depth maps. We detail the different components of our generative prior in Sec 3.2 and the subsequent 3DGS optimization in Sec 3.4. We end the section with a final pose alignment step for a given set of test views in Sec 3.5.
@@ 3.1 Algorithm Overview
Algorithm 1 GScenes Algorithm
1:Sparse input image set =$\mathcal{I}=\{I_{1},I_{2},\dots I_{M}\}$
2:Set of 3D Gaussians $\mathcal{G}$ and camera poses $\pi=\{\pi_{1},\pi_{2},\dots\pi_{M}\}$
$\hat{\mathcal{I}}\leftarrow\mathcal{I}$
4:$\mathcal{G}\leftarrow$ Optimize 3DGS with DUSt3r point cloud for $1k$ iterations
5:for $N$ iterations do
6:     $\pi\leftarrow$ Sample novel camera pose.
7:     $\mathbf{I},\mathbf{D}\leftarrow R_{\pi}(\mathcal{G})$ Render from camera $\pi$
8:     $\mathbf{I}\leftarrow$ Refine$(\mathbf{I})$
9:     $\hat{\mathbf{I}}\leftarrow\ \hat{\mathbf{I}}\cup\{\mathbf{I}\}$
10:     $\mathcal{G}\leftarrow$ Optimize 3DGS for $k$ iterations
11:end for
An overview of our method is given in Fig 1 and Alg 1. We initialize GScenes with an incomplete dense Gaussian point cloud reconstruction from sparse input images using InstantSplat Fan et al. (2024). We use this incomplete Gaussian scene representation as an implicit geometric prior and sample novel views along a smooth B-spline trajectory fitted to training views. We use our customized RGBD diffusion prior to generate plausible novel views. In addition to CLIP features of source images for context, we devise a novel 3DGS confidence measure to effectively guide our diffusion model towards empty regions and potential artifacts in a novel view render. We then run $10k$ iterations of 3DGS optimization, sampling a novel view before each densification step (Fu et al., 2023) to obtain our final scene representation.
@@
The DUSt3r pipeline gives us a pixel-aligned dense-stereo point cloud $\mathbf{P}\in\mathbb{R}^{S\times 3}$, camera intrinsics $\{\mathbf{K_{i}}\in\mathbb{R}^{3\times 3}\}_{i=1}^{M}$ and extrinsics $\{\mathbf{E_{i}}=[\mathbf{R_{i}|T_{i}}]\}_{i=1}^{M}$ for our $M$ input images. Nevertheless, both $\mathbf{P}$ and the estimated poses demonstrate sub-optimal alignment compared to those generated by COLMAP (Schönberger and Frahm, 2016) from the dense observation dataset. Consequently, similar to the approach used in InstantSplat, we start by initializing 3D Gaussians at each location within the globally aligned point cloud $\mathbf{P}$. Subsequently, we jointly optimize both the Gaussian attributes and camera parameters over the course of $1k$ iterations without incorporating any form of Adaptive Density Control.
3.2 RGBD Diffusion Priors for NVS
Despite the initial geometric prior, many areas of the scene remain unobserved due to the sparsity of the input set, leading to areas of no Gaussians (0-Gaussians) in the 3D representation $\mathcal{G}$ obtained from InstantSplat and zero opacity regions, floaters, and artifacts in novel views. Regularization techniques or 3D foundational priors cannot help with inferring missing details, prompting us to develop an instruction-following RGBD diffusion model comprising a variational autonecoder with encoder $\mathcal{E}$, decoder $\mathcal{D}$ and UNet $\bm{\epsilon_{\theta}}$ operating in the latent space of the VAE. $\bm{\epsilon_{\theta}}$ learns to predict noise in a noisy latent $z_{t}$ conditioned on a RGBD image $\hat{I}$, corresponding confidence map $\mathcal{C}$, clip conditioning $c_{clip}$, and a text instruction $y$. $\hat{I}$ and $\mathcal{C}$ are obtained by rendering $\mathcal{G}$ from novel views.
We initialize $\bm{\epsilon_{\theta}}$ with the pretrained weights of the Stable-Diffusion-2 checkpoint(Rombach et al., 2022) to leverage its image generation capabilities for RGBD image-to-image synthesis. We introduce additional channels in the first convolutional layer of $\bm{\epsilon_{\theta}}$ to support image conditioning with $\hat{I}$. Specifically, we concatenate $z_{t}$, $\mathcal{E}(\hat{I})$ and $\mathcal{\hat{C}}$, where $\mathcal{\hat{C}}$ is $\mathcal{C}$ downsampled to the resolution of the latent space. Weights of all newly added channels are initialized to zero. However, conditioning on $\hat{I}$ and $\mathcal{C}$ alone does not ensure novel view synthesis coherent with the observed inputs. In the absence of explicit 3D conditioning, we augment our generative model with scene context using CLIP features. We extract CLIP features $c_{clip}\in\mathcal{R}^{M\times d}$ of the $M$ training images to modulate intermediate feature layers in $\bm{\epsilon_{\theta}}$ However, instead of the usual cross-attention mechanism, we resort to a more parameter-efficient approach by using Feature-wise Linear Modulation (FiLM) layers (Perez et al., 2018). These layers work by generating scaling and shifting parameters for the UNet feature maps, allowing $c_{clip}$ to modulate intermediate activations in the UNet. We first pass $c_{clip}$ through trainable self-attention modules to capture shared semantic information across views and create a unified, context-aware representation.
$c_{attn}=\text{SelfAttention}(c_{clip})$ (1)
$c_{attn}\in\mathcal{R}^{d}$ is next passed through layer-specific fully connected layers to predict scale and shift parameters $\gamma,\beta$ as:
$\gamma^{(l)},\beta^{(l)}=FC^{(l)}(c_{attn})$ (2)
Let $\mathbf{F}^{(l)}\in\mathcal{R}^{k\times h\times w}$ represent a feature map in $\bm{\epsilon_{\theta}}$ at a layer $l$, where $k$ is the number of channels, $h$ and $w$ are the spatial dimensions. The FiLM modulation is applied element-wise as:
$\mathbf{F}_{mod}^{(l)}=\text{FiLM}(\mathbf{F}^{(l)}|\gamma^{(l)},\beta^{(l)})=\gamma^{(l)}\cdot\mathbf{F}^{(l)}+\beta^{(l)}$ (3)
This conditioning enables the dynamic adaptation of the UNet’s feature maps facilitated by scale and shift parameters conditioned on the CLIP embedding, thereby directing the UNet’s output to align with the semantic features captured by CLIP. We implement FiLM modulation with CLIP embeddings exclusively within the down and mid blocks of the UNet, as this approach allows for the effective integration of conditional information without perturbing the generation process in the up(sample) blocks.

### v2

@@ Abstract
In this work, we introduce a generative approach for pose-free (without camera parameters) reconstruction of $360^{\circ}$scenes from a sparse set of 2D images. Pose-free scene reconstruction from incomplete, pose-free observations is usually regularized with depth estimation or 3D foundational priors. While recent advances have enabled sparse-view reconstruction of large complex scenes (with high degree of foreground and background detail) with known camera poses using view-conditioned generative priors, these methods cannot be directly adapted for the pose-free setting when ground-truth poses are not available during evaluation. To address this, we propose an image-to-image generative model designed to inpaint missing details and remove artifacts in novel view renders and depth maps of a 3D scene. We introduce context and geometry conditioning using Feature-wise Linear Modulation (FiLM) modulation layers as a lightweight alternative to cross-attention and also propose a novel confidence measure for 3D Gaussian splat representations to allow for better detection of these artifacts. By progressively integrating these novel views in a Gaussian-SLAM-inspired process, we achieve a multi-view-consistent 3D representation. Evaluations on the MipNeRF360 and DL3DV-10K benchmark dataset demonstrate that our method surpasses existing pose-free techniques and performs competitively with state-of-the-art posed (precomputed camera parameters are given) reconstruction methods in complex $360^{\circ}$scenes.
Figure 1: Given sparse pose-free images as input, GScenes reconstructs a 3D scene in 5 minutes by iteratively fusing novel view renders and depth maps with an underlying 3D Gaussian representation. Typical pose-free baselines built with geometric priors struggle with reconstructing $360^{\circ}$scenes from sparse inputs due to the absence of generative priors. GScenes comprises a latent diffusion model capable of inpainting missing details and removing Gaussian artifacts in novel view renders, thereby enabling generation of full $360^{\circ}$scenes.
@@ 1 Introduction
Reconstructing high-quality 3D scenes from sparse images remains a fundamental challenge in computer vision. While recent methods employ various priors to stabilize NeRFs (Mildenhall et al., 2020) or Gaussian splats (3DGS) (Kerbl et al., 2023) in under-constrained scenarios, they typically require accurate camera parameters derived from dense observations—a restrictive assumption for real-world applications. Pose estimation from sparse views is inherently challenging; both traditional Structure from Motion and recent foundational models (Wang et al., 2024; Leroy et al., 2024) struggle with insufficient matching features. Current pose-free 3DGS approaches integrate monocular depth (Ranftl et al., 2020), semantic segmentation (Kirillov et al., 2023), or 3D priors (Wang et al., 2024), but fail on complex $360^{\circ}$scenes with sparse coverage, highlighting the need for additional generative regularization.
Although most previous sparse view 3D scene reconstruction works are posed and require ground-truth camera poses (DSNeRF (Deng et al., 2022), DNGaussian (Li et al., 2024), SparseGS (Xiong et al., 2023), SparseNeRF (Wang et al., 2023a), DietNeRF (Jain et al., 2021), FreeNeRF (Yang et al., 2023), RegNeRF (Niemeyer et al., 2022), DiffusioNeRF (Wynn & Turmukhambetov, 2023)), recent works like COGS (Jiang et al., 2024) and InstantSplat (Fan et al., 2024) have found initial success in the pose-free sparse view reconstruction problem. Although they have been successful in reducing 3D reconstruction-related artifacts such as blur, floaters, color, and streak-like artifacts, they lack generative capabilities for a complete $360^{\circ}$reconstruction, which requires robust priors from powerful generative models (Saharia et al., 2022; Rombach et al., 2022). In addition, most of these methods lack mechanisms to robustly identify and locate these artifacts in the reconstructed scenes. Recent methods like ZeroNVS (Sargent et al., 2024), Reconfusion (Wu et al., 2024), and CAT3D (Gao* et al., 2024) incorporate 3D view conditioning for realistic extrapolation but depend on accurate poses and cannot be trivially extended to the pose-free problem. Gaussian Object (Yang et al., 2024), iFusion (Wu et al., 2023), and UpFusion (Nagoor Kani et al., 2024) do provide pose-free generative solutions but remain limited to only object reconstruction rather than scene reconstruction. Other generative solutions like ZeroNVS, ReconFusion, and CAT3D use stronger priors over regularization-based techniques like FreeNeRF, RegNeRF, and DietNeRF but rely on large-scale 3D datasets, video datasets, and compute resources not available to the average researcher. Furthermore, works like ReconFusion and CAT3D remain closed-source with no access to their models, data or other reproducibility data, which limits their usage for the community. We summarize such existing methods and their use cases and attributes in Fig 2.
Method Pose-free Open-source Generative Priors Scene Reconstruction
FreeNeRF, DietNeRF,
RegNeRF, DN-Gaussian,
SparseGS, SparseNeRF
DiffusioNeRF, ZeroNVS
ReconFusion, CAT3D
Gaussian Object,
iFusion, UpFusion
InstantSplat, COGS
GScenes (Ours)
Figure 2: Comparison of sparse-view reconstruction methods. Methods are grouped based on their requirement for accurate camera poses, open-source availability, need for generative priors, and applicability to large-scale scene reconstruction.
@@
To alleviate these challenges, we present GScenes, an efficient approach using 3D foundational and RGBD (RGB image and depth) generative priors for pose-free sparse-view reconstruction of complex $360^{\circ}$scenes. We first estimate a point cloud and approximate camera parameters using MASt3R (Leroy et al., 2024), then jointly optimize Gaussians and cameras with 3DGS. Novel views generated from an elliptical trajectory fitted to the training views contain artifacts that our diffusion prior refines to further optimize the scene. We condition a Stable Diffusion UNet (Rombach et al., 2022) on estimated cameras, context, 3DGS renders, depth maps, and a confidence map capturing artifacts and missing details. Our proposed confidence measure is designed by combining per-pixel light transmittance with Gaussian density, which provides a reliable conditioning signal to the Stable Diffusion UNet on the presence of empty regions and Gaussian artifacts in novel view renders and depth maps. For training our diffusion prior, we create our own dataset of 171, 461 samples using scenes from open-source multiview datasets. Each sample is a pair of clean RGBD images and images with empty regions and Gaussian artifacts, along with other conditioning signals (confidence map, clip features, etc). Augmented variants of both the Stable Diffusion VAE and UNet are finetuned using this synthetic dataset. Despite using weaker and cheaper generative priors than pose-dependent methods, we demonstrate competitive performance against closed source methods like ReconFusion (Wu et al., 2024) and CAT3D (Gao* et al., 2024) while outperforming other techniques without requiring million-scale multi-view or video data or extensive compute resources. Our contributions include:
  • An image-to-image RGBD generative model for synthesizing plausible novel views from sparse pose-free images, using lightweight FiLM modulation (Perez et al., 2018) instead of cross-attention
  • A confidence measure to detect artifacts in novel view renders, guiding our diffusion model toward effective novel view synthesis
  • Integration of diffusion priors with MASt3R’s geometry prior, enabling efficient scene reconstruction previously requiring 3D-aware video diffusion
  • Superior performance compared to recent regularization and generative prior-based open-source methods for 3D scene reconstruction
  • An open-source low-cost solution with lower data and compute requirements compared to state-of-the-art posed reconstruction methods.
See Appendix for further discussion and Related Works (Appendix B).
2 Method
@@
This section begins with an overview of our method in Sec 2.1, detailing our approach for reconstructing a 3D scene from a sparse set of uncalibrated 2D images. In Sec 2.1, we describe how we initialize a Gaussian point cloud using MASt3R and 3DGS to provide a coarse 3D representation. Sec 2.2 introduces our RGBD image-to-image generative model, which refines rendered novel views by correcting artifacts and filling in missing details. In Sec 2.3, we propose a confidence measure based on cumulative transmittance and Gaussian density to guide the generative diffusion model in identifying unreliable regions in novel-view renders. Sec 2.4 outlines our synthetic dataset creation process, which enables training the generative model with high-quality RGBD supervision. Sec 2.5 details our depth-augmented autoencoder finetuning process to improve latent-space encoding for RGBD data. Sec 2.5.1 explains the fine-tuning of our UNet with synthesized training data to generate photorealistic and geometrically consistent novel views. In Sec 2.5.2, we describe the inference process of our generative model, where novel views are synthesized using RGBD renders and confidence maps. Sec 2.6 presents our iterative 3D Gaussian optimization strategy that progressively integrates novel view constraints into the scene representation. Finally, Sec 2.7 describes a test-time pose alignment step that refines camera poses to align the rendered image with a given test view before evaluation.
@@ 2.1 Algorithm Overview
Algorithm 1 Gaussian Scenes Training
1:Sparse image set =$\mathcal{I}=\{I_{1},\dots,I_{M}\}$, densification interval $k$, iterations $N$
2:Initialize 3D Gaussian primitives $\mathcal{G}$ from MASt3R 3D point cloud $\mathbf{P}$
3:Optimize $\mathcal{G}$ using 3DGS on input views $\mathcal{I}$ and estimated cameras $\pi_{train}=\{\pi_{1},\dots,\pi_{M}\}$ for $1k$ iterations
$\hat{\mathcal{I}}\leftarrow\mathcal{I}$triangleright Initialize training images
5:for $t=1,\dots,N$ do
6:     if $t\mod k=0$ then
7:         Sample novel pose $\pi_{novel}$ from elliptical trajectory fitted to $\pi_{train}$
8:         $(\hat{I},\hat{D})\leftarrow R_{\pi_{novel}}(\mathcal{G})$triangleright Render image and depth map
9:         Compute confidence map $\mathcal{C}$ (Eq. 2)
10:         Extract semantic context & geometry embeddings $\{c_{clip},c_{geo}\}$triangleright Sec 2.2.2
11:         $(I_{ref},D_{ref})\leftarrow\text{Refine}(\hat{I},\hat{D},\mathcal{C},c_{clip},c_{geo})$triangleright Refine $\hat{I}$, $\hat{D}$ with generative priors (Sec 2.5.2)
12:         ^^$\hat{\mathcal{I}}\leftarrow\hat{\mathcal{I}}\cup\{I_{ref}\}$triangleright Synthesized novel view added as training sample
13:     end if
14:     Sample pose $\pi_{i}$ from $\pi_{train}$
15:     Render $(I_{\pi},D_{\pi})\leftarrow R_{\pi_{i}}(\mathcal{G})$
16:     Optimize 3D Gaussians $\mathcal{G}$triangleright Eq. 6 or 3DGS loss
17:end for
Figure 4: Camera Trajectory Visualization for Novel View Synthesis in pose-free sparse-view setting.
An overview of our method is given in Fig 3 and Alg 1. We initialize GScenes with an incomplete dense Gaussian point cloud reconstruction from sparse input images using MASt3R and 1k iterations of 3DGS. Note that InstantSplat proposes the same framework for sparse-view reconstruction, but with a DUSt3R initialization. We choose MASt3R to initialize scene geometry instead due to its superior performance in the sparse-view setting. We use this incomplete scene representation as an implicit geometric prior and sample novel views along a smooth elliptical trajectory fitted to training views. An example trajectory is shown in Fig 4. We then use our RGBD generative prior to synthesize plausible novel views. In addition to CLIP features of source images for context and plücker embeddings of source and target cameras for geometric conditioning, we devise a novel 3DGS confidence measure to effectively guide our generative model towards empty regions and potential artifacts in a novel view render. We then run $10k$ iterations of 3DGS optimization, sampling a novel view before each densification step (Fu et al., 2023) to obtain our final scene representation.
@@
The MASt3r initialization gives us a pixel-aligned dense-stereo point cloud $\mathbf{P}\in\mathbb{R}^{S\times 3}$, camera intrinsics $\{\mathbf{K_{i}}\in\mathbb{R}^{3\times 3}\}_{i=1}^{M}$ and extrinsics $\{\mathbf{E_{i}}=[\mathbf{R_{i}|T_{i}}]\}_{i=1}^{M}$ for our $M$ input images. Nevertheless, both $\mathbf{P}$ and the estimated poses demonstrate sub-optimal alignment compared to those generated by COLMAP from the dense observation dataset. We cannot also use COLMAP for camera or geometry estimation as the method fails when there is minimal overlap between training images, as in our setting. Consequently, similar to the approach in InstantSplat, we initialize 3D Gaussians at each location in the globally aligned point cloud $\mathbf{P}$. We then jointly optimize both the Gaussian attributes and camera parameters over the $1k$ iterations. The MASt3R point cloud is already quite dense ( 1.5-3M points), which alleviates the need for any form of Adaptive Density Control (Kerbl et al., 2023).
2.2 RGBD Generative Priors for Novel View Synthesis (NVS)
Reconstructing complete 3D scenes from sparse observations requires inferring content in unobserved regions—a fundamental challenge that geometric regularization and 3D priors alone cannot adequately address. We introduce a diffusion-based generative approach that leverages 2D image priors to synthesize plausible content.
2.2.1 Generative Model Architecture
Despite our initial geometric reconstruction, sparse input views inevitably result in regions with no Gaussian primitives (“0-Gaussians”), causing empty areas and artifacts in novel views. Unlike regularization-based methods that merely constrain optimization without generating content, our approach directly synthesizes missing scene details.
Our model comprises of (1) A variational autoencoder (encoder $\mathcal{E}$, decoder $\mathcal{D}$) operating in a compressed latent space. (2) A UNet denoiser $\epsilon_{\theta}$ predicting noise in diffused latent $z_{t}$ (3) A multi-modal conditioning incorporating RGBD renders, confidence maps, semantic context, and geometric information.
The UNet $\epsilon_{\theta}$ receives four inputs: an artifact-laden RGBD image $\hat{I}$, a confidence map $\mathcal{C}$ identifying unreliable regions, CLIP features $c_{\text{clip}}$ of $\mathcal{I}$ providing semantic context, and camera encodings $c_{\text{geo}}$ establishing geometric relationships between views.
2.2.2 Multi-modal Conditioning
We initialize $\epsilon_{\theta}$ with pretrained Stable-Diffusion-2 model weights and expand the first convolutional layer to accept additional inputs by concatenating the noisy latent $z_{t}$, the encoded RGBD image $\mathcal{E}(\hat{I})$, the confidence-weighted encoded image $\mathcal{E}(\hat{I}\cdot\mathcal{C})$ and the downsampled confidence map $\hat{\mathcal{C}}$. To ensure view coherence and geometric consistency, we incorporate following conditioning signals:
Semantic Context: CLIP features $c_{\text{clip}}\in\mathbb{R}^{M\times d}$ from source images serve as semantic anchors, ensuring generated content remains consistent with observed scene elements.
Geometric Information: For each camera with center $\mathbf{o}$ and forward axis $\mathbf{d}$, we compute its plücker coordinates $\mathbf{r}=(\mathbf{d},\mathbf{o}\times\mathbf{d})\in\mathbb{R}^{6}$ and apply frequency encoding for obtaining higher-dimensional features:
$\mathbf{r}\mapsto[\mathbf{r},\sin(f_{1}\pi\mathbf{r}),\cos(f_{1}\pi\mathbf{r}),\cdots,\sin(f_{K}\pi\mathbf{r}),\cos(f_{K}\pi\mathbf{r})]$ (1)
where $K=6$ is the number of Fourier bands, and $f_{k}$ are equally spaced frequencies. This yields a 78-dimensional embedding for each camera ($c_{\text{geo}}\in\mathbb{R}^{(M+1)\times 78}$), capturing geometric relationships between viewpoints. Plücker coordinates were originally introduced by LFNs (Sitzmann et al., 2021) for per-pixel parameterization of a ray. We instead obtain a single representation per camera using extrinsic parameters $\mathbf{E_{i}}$ for obtaining $\mathbf{o}$ and $\mathbf{d}$.
2.2.3 Parameter-Efficient FiLM Conditioning
MASt3R + 3DGS Diffusion Sample (UpBlock Conditioning) Diffusion Sample (Ours)
3 views Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Incorporating context and geometry conditioning in the up blocks of the UNet negatively impacts latent and subsequent image reconstruction.
We employ Feature-wise Linear Modulation (FiLM) instead of cross-attention for incorporating context and geometry information, achieving both computational efficiency and strong performance. We process context and geometry embeddings through self-attention to capture inter-view relationships as: $c^{i}_{\text{attn}}=\text{SelfAttention}(c_{i});\quad i\in\{\text{clip},\text{geo}.\}$ We then generate scaling and shifting parameters via layer-specific networks as: $\gamma^{(l)},\beta^{(l)}=\text{FC}^{(l)}(c^{i}_{\text{attn}}).$ Finally, we modulate the UNet feature maps through element-wise operations as: $\mathbf{F}_{\text{mod}}^{(l)}=\gamma^{(l)}\cdot\mathbf{F}^{(l)}+\beta^{(l)}.$
Critically, we apply FiLM modulation only to down and mid blocks of the UNet—not up blocks—based on empirical evidence showing this selective application yields optimal results (Fig. 5).
Our FiLM-based approach requires only 8.14M parameters (7.38M for CLIP features, 758K for pose embeddings) compared to 29.8M for an equivalent cross-attention implementation—a 3$\times$ reduction while maintaining comparable quality, enabling more efficient training and faster inference during iterative reconstruction.